OpenStudy (anonymous):

A new 13.3'' Macbook Pro costs $1,200.00, and a one year old 13.3'' Macbook Pro costs $1,000.00. What is the percentage it decreased in one year?

OpenStudy (ash2326):

New Macbook Cost $1,200.00 One year old Macbook cost $1,000.00 Percentage decrease \[\frac{New-Old}{New} \times 100=\frac{1200-1000}{1200} \times 100\] Can you finish this?

OpenStudy (anonymous):

is the answer 16%?

OpenStudy (ash2326):

Yeah it's 16.66%

OpenStudy (anonymous):

so I can just write it as 16%

OpenStudy (ash2326):

No, that'd be a totally different answer. Write it as 16.66 or 16.7